3x^4-2x^2-1=0 find the real solution of this equation
Answer by lwsshak3(11628)

3x^4-2x^2-1=0 find the real solution of this equation
...
3x^4-2x^2-1=0
let u=x^2
u^2=x^4
..
3u^2-2u-1=0
(u-1)(3u+1)=0
u=1
or
u=-1/3
..
x^2=u=-1/3
x=±√(-1/3) (2 imaginary solutions)
x^2=u=1
x=±√1=±1 (2 real solutions)